# Fractional Semilinear Equations on Hyperbolic Spaces

### v1

@@ Abstract.
We study a semilinear equation involving the fractional Laplacian on the hyperbolic space $\mathbb{H}^{n}$. Unlike in conformally compact Einstein manifolds, the fractional Laplacian on $\mathbb{H}^{n}$ does not enjoy conformal covariance. By employing Helgason-Fourier analysis, we explicitly derive the Green’s function of the fractional Laplacian on $\mathbb{H}^{n}$ and study its asymptotic behaviors. We then apply a direct method of moving planes to the integral form of the equation, establishing symmetry of solutions and nonexistence of positive solutions in the critical and subcritical cases, respectively. In addition, we develop several maximum principles on hyperbolic space.
@@ 1. Introduction
We point out here that the moving plane method on $\mathbb{H}^{n}$ was first seen in the work of Kumaresan-Prajapat [33, 34], where they established the analog result of Gidas-Ni-Nireneberg type, as well as solved an overdetermined problem on $\mathbb{H}^{n}$, which is an analogue of the problem on $\mathbb{R}^{n}$ initiated by Serrin [52] (see also [27], [49]). The hyperbolic space serves as one of the most important models of Riemannian manifolds with constant curvature. For definitions and basic properties of hyperbolic spaces, we refer to Section 2. The axial symmetry of solutions to a class of integral equations on half spaces was studied by Lu-Zhu in [50]. More recently, the overdetermined problem of fractional order equations on hyperbolic spaces was studied by Li-Lu-Wang [38]. The concept of the moving plane method in [33, 34] was further developed in the work of Almeida-Ge [4] and Almeida-Damascelli-Ge [3], where they took advantage of the foliation structure of $\mathbb{H}^{n}$ (for details see Section 2). Li, Lu and Yang [40] obtained a higher order symmetry result of the Brezis-Nirenberg problem on the hyperbolic spaces [40]. Recently. with Li and Lu, the author studied general higher order equations [37] by using the moving plane method, and later obtained the classification result using the moving sphere method [39] on the hyperbolic space.
@@
$(-\Delta_{\mathbb{H}^{n}})^{s}u=u^{p}(x),\quad x\in\mathbb{H}^{n}.$ (1.7)
We say that a nonnegative function $u\in H^{s}(\mathbb{H}^{n})$ is a weak solution of (1.7) if for any test function $\varphi\in C_{c}^{\infty}(\mathbb{H}^{n})$, there holds
@@
Let $n\geqslant 2$, $0<s<1$. Suppose that $u\in L^{q}(\mathbb{H}^{n}),q>p$ is a nonnegative weak solution of (1.7). Then
  1. (1)
    If $1<p<\frac{n+2s}{n-2s}$ and $u\in L^{p+1}(\mathbb{H}^{n})$ , then $u\equiv 0$.
  2. (2)
    If $p=\frac{n+2s}{n-2s}$ and $u\in L^{q}$, where $q>p$, then there exists a point $x_{0}\in\mathbb{H}^{n}$ such that $u$ is radially symmetric and nondecreasing about $x_{0}$.
@@
This equivalence plays an essential role in applying the moving plane method to prove Theorem 1.1. The Helgason-Fourier analysis on general symmetric spaces was first developed by Helgason [30, 31]. It has been used in harmonic analysis and PDEs; see e.g. [5, 7] and the references therein. In addition, the Helgason-Fourier analysis plays a key role in the establishment of higher order Hardy-Sobolev-Maz’ya’s inequalities (see [45, 46, 47, 20]). Using Fourier transform, we are able to derive the precise expression of the Green’s function $G_{s}(\rho)$ and study its asymptotic behaviors near zero and infinity, which are the key ingredients in the moving plane method.
@@
For $s\in(0,1)\setminus\mathbb{N}$, there holds that
@@ 2.2. Hardy-Littlewood-Sobolev inequality
$\|I_{\alpha}f\|_{L^{q}(\mathbb{H}^{n})}\leqslant C_{n,s,}\,\|f\|_{L^{p}(\mathbb{H}^{n})}.$ (2.2)
@@ 2.3. Sobolev inequality for fractional Laplacians on $\mathbb{H}^{n}$
For a general survey and the manifold setting, one may be referred to Hebey [29] and Lions [54, 55] for concentration-compactness techniques. For semilinear PDEs exploiting radial compactness in hyperbolic space see Mancini-Sandeep [51] and related works.
@@ 2.4. The Helgason-Fourier transform on hyperbolic spaces
the generalized eigenfunctions of the Laplace-Beltrami operator $\Delta_{\mathbb{H}^{n}}$ on the ball model $\mathbb{B}^{n}$ with eigenvalue $(\lambda^{2}+\frac{(n-1)^{2}}{4})$. The Fourier transform of a function $f$ on $\mathbb{H}^{n}$ is defined as
@@
$f(x)=D_{n}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\hat{f}(\lambda,\zeta)e_{\lambda,\zeta}(x)|\mathfrak{c}(\lambda)|^{-2}d\lambda d\sigma,$
where $D_{n}=(2^{3-n}\pi|\mathbb{S}^{n-1}|)^{-1}$ and $\mathfrak{c}(\lambda)$ is the Harish-Chandra’s $\mathfrak{c}$-function given by
$\mathfrak{c}(\lambda)=\frac{2^{n-1-i\lambda}\Gamma(n/2)\Gamma(i\lambda)}{\Gamma(\frac{n-1+i\lambda}{2})\Gamma(\frac{1+i\lambda}{2})}.$
@@
$\int_{\mathbb{B}^{n}}|f(x)|^{2}dV=D_{n}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\hat{f}(\lambda,\zeta)|\mathfrak{c}(\lambda)|^{-2}d\lambda d\sigma.$
@@ 2.5. Foliations of hyperbolic spaces
Let $U=\mathbb{H}^{n}\cap\{x_{1}=0\}$ and $U_{t}=A_{t}(U)$, $\mathbb{H}^{n}$ is then foliated by $U_{t}$ and $\mathbb{H}^{n}=\bigcup_{t\in\mathbb{R}^{n}}U_{t}$. The reflection $I$ is an isometry such that $I^{2}=Id$ and $I$ fixes the hypersurface $U$, by $I(x_{0},x_{1},x_{2},\cdots,x_{n})=(x_{0},-x_{1},x_{2},\cdots,x_{n})$. Moreover, the reflecton w.r.t. $U_{t}$ is defined as $I_{t}=A_{t}\circ I\circ A_{-t}$, and $U_{t}$ is fixed by $I_{t}$.
@@ 3. Fractional Laplacian on hyperbolic spaces
We point out here that the above two definitions of fractional Laplacians are not equivalent. Although the later is also defined via an extension problem, the manifold $\mathbb{H}^{n}\times(0,\infty)$ is not conformally compact Einstein. In fact, the former operator lives on the conformal infinite of $X$ while the latter is defined on the bulk manifold $X$ itself. As a consequence, the fractional Laplacian we treat here does not coincide with the scattering operator on hyperbolic spaces, which is not conformally covariant.
@@
$L_{s}(x,y)=\int_{\mathbb{S}^{n}}\bar{h}_{\lambda,\theta}(x)h_{\lambda,\theta}(y)\,d\theta.$ (3.2)
We have the following explicit formulae of $L_{s}(x,y)$ given by Banica [6].
@@
$L_{s}(x,y)=C_{n}\,(\sinh\rho)^{-\frac{n-2}{2}}P_{-\frac{1}{2}+is}^{-\frac{n-2}{2}}(\cosh\rho),$ (3.3)
@@
$L^{b}_{s}(\rho)=C_{n}\left(-\frac{1}{\sinh\rho}\frac{d}{d\rho}\right)^{\frac{n-1}{2}}\cos(\lambda\rho),$ (3.4)
@@
$L^{b}_{s}(\rho)=C_{n}\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(-\frac{1}{\sinh r}\frac{d}{dr}\right)^{\frac{n}{2}}\cos(\lambda r)\,dr.$ (3.5)
The kernel of Green’s function $G_{s}(x,y)$ can be considered as the inverse of the fractional Laplacian. In [7], the authors provided the information of $\mathcal{K}_{s}$, the compuation for $G_{s}$ is similar, and we present the detailed proof here for completeness.
@@
$\Big(\frac{\partial_{\rho}}{\sinh\rho}\Big)^{m}\sim 2^{m}e^{-m\rho}\partial_{\rho}^{m}-(2^{m+1}e^{-m\rho}+2^{m}e^{-\rho})\partial_{\rho}^{m-1},\qquad\rho\to\infty.$
Finally, we prove the following monotonicity property of the kernel $G_{n,s}(\rho)$.
@@
If $n=2m+1$, we have
$G_{2m+1,s}(\rho)=C_{1}\left(-\frac{\partial_{\rho}}{\sinh\rho}\right)^{m}\left(\rho^{-\frac{1-2s}{2}}K_{s-\frac{1}{2}}\left(\frac{n-1}{2}\rho\right)\right).$
Thus,
$\frac{d}{d\rho}G_{2m+1,s}(\rho)=C_{1}(-\sinh\rho)\left(-\frac{\partial_{\rho}}{\sinh\rho}\right)^{m+1}\left(\rho^{-\frac{1-2s}{2}}K_{s-\frac{1}{2}}\left(\frac{n-1}{2}\rho\right)\right)$
$=C_{1}(-\sinh\rho)\cdot G_{2m+3,s}(\rho)<0.$
If $n=2m$, we have
$G_{2m,s}(\rho)$ $=C_{1}\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\left(-\frac{\partial_{r}}{\sinh r}\right)^{m}\left(r^{-\frac{1-2s}{2}}K_{s-\frac{1}{2}}\left(\tfrac{n-1}{2}r\right)\right)dr.$
Now if we set $b=\sqrt{\cosh r-\cosh\rho}$, we get
$\frac{d}{d\rho}G_{2m,s}(r)=\frac{1}{\sqrt{\pi}}\int_{0}^{\infty}\frac{d}{d\rho}G_{2m+1,s}(r(b,\rho))db,$
where $r(b,\rho)$ is defined by $b=\sqrt{\cosh r-\cosh\rho}$. Since
$\frac{d}{d\rho}G_{2m+1,s}(r)=\frac{dG_{2m+1,s}(r)}{dr}\frac{dr}{d\rho}=\frac{dG_{2m+1,s}(r)}{dr}\frac{\sinh\rho}{\sinh r}<0,$
we have $\frac{d}{d\rho}G_{2m,s}(r)<0$. The proof is now concluded. ∎
@@
Proposition 4.1.
Let $s\in(0,1),u\in L^{q}_{\mathrm{loc}}(\mathbb{H}^{n}),q>p$. Then the weak solution to
$(-\Delta_{\mathbb{H}^{n}})^{s}u=u^{p}$ (4.1)
satisfies the integral equation
$u(x)=\int_{\mathbb{H}^{n}}\mathcal{K}_{s}(\rho(x,y))u(y)^{p}dV_{y},$ (4.2)
where $\mathcal{K}_{s}(\rho)$ is the fractional kernel on $\mathbb{H}^{n}$ defined in Section (1.6). Conversely, if $u$ solves the integral equation (4.2), then $u$ is a solution of the (4.1).
@@
$(-\Delta_{\mathbb{H}^{n}})^{s}\varphi(x)=\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\Bigl(\lambda^{2}+\frac{(n-1)^{2}}{4}\Bigr)^{s}\widehat{\varphi}(\lambda,\theta)\,e_{\lambda,\theta}(x)\,|c(\lambda)|^{-2}\,d\theta\,d\lambda.$ (4.3)
@@
$\int_{\mathbb{H}^{n}}u(x)(-\Delta_{\mathbb{H}^{n}})^{s}\varphi(x)\,dV_{x}$ $=\int_{\mathbb{H}^{n}}u(x)\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\Bigl(\lambda^{2}+\tfrac{(n-1)^{2}}{4}\Bigr)^{s}\widehat{\varphi}(\lambda,\theta)e_{\lambda,\theta}(x)|c(\lambda)|^{-2}\,d\theta\,d\lambda\,dV_{x}$
$=\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\Bigl(\lambda^{2}+\tfrac{(n-1)^{2}}{4}\Bigr)^{s}\widehat{\varphi}(\lambda,\theta)\left[\int_{\mathbb{H}^{n}}u(x)e_{\lambda,\theta}(x)\,dV_{x}\right]|c(\lambda)|^{-2}\,d\theta\,d\lambda.$
Similarly, the right side becomes
$\int_{\mathbb{H}^{n}}u(x)^{p}\varphi(x)\,dV_{x}$ $=\int_{\mathbb{H}^{n}}u(x)^{p}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\widehat{\varphi}(\lambda,\theta)e_{\lambda,\theta}(x)|c(\lambda)|^{-2}\,d\theta\,d\lambda\,dV_{x}$
$=\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\widehat{\varphi}(\lambda,\theta)\left[\int_{\mathbb{H}^{n}}u(x)^{p}e_{\lambda,\theta}(x)\,dV_{x}\right]|c(\lambda)|^{-2}\,d\theta\,d\lambda.$
@@
$\Bigl(\lambda^{2}+\tfrac{(n-1)^{2}}{4}\Bigr)^{s}\int_{\mathbb{H}^{n}}u(x)e_{\lambda,\theta}(x)\,dV_{x}=\int_{\mathbb{H}^{n}}u(x)e_{\lambda,\theta}(x)\,dV_{x}$
Thus, by inversion formula, we obtain
$u(x)$ $=\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\left(\lambda^{2}+\tfrac{(n-1)^{2}}{4}\right)^{-s}\widehat{u^{p}}(\lambda,\theta)h_{\lambda,\theta}(x)|c(\lambda)|^{-2}d\theta d\lambda$
$=\int_{\mathbb{H}^{n}}u(y)^{p}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\left(\lambda^{2}+\tfrac{(n-1)^{2}}{4}\right)^{-s}\bar{h}_{\lambda,\theta}(x)h_{\lambda,\theta}(y)|c(\lambda)|^{-2}d\theta d\lambda dV_{y}$
@@
Here $G_{s}(\rho)$ is the Green’s function of the fractional Laplacian on $\mathbb{H}^{n}$, given by
$G_{s}(\rho)=\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\left(\lambda^{2}+\tfrac{(n-1)^{2}}{4}\right)^{-s}\bar{h}_{\lambda,\theta}(x)h_{\lambda,\theta}(y)|c(\lambda)|^{-2}d\theta d\lambda.$
The reversed direction follows from similar calculations and we omit the details. ∎

### v2

# Symmetry of Solutions to Fractional Semilinear Equations on Hyperbolic Spaces

@@ Abstract.
We study a semilinear equation involving the fractional Laplacian on the hyperbolic space $\mathbb{H}^{n}$. Unlike in conformally compact Einstein manifolds, the fractional Laplacian on $\mathbb{H}^{n}$ does not enjoy conformal covariance. By employing Helgason-Fourier analysis, we explicitly derive the Green’s function of the fractional Laplacian on $\mathbb{H}^{n}$ as well as its asymptotic behaviors. We then apply a direct method of moving planes to the integral form of the equation, and show that nonnegative weak solutions are symmetric. In addition, we extend several maximum principles to hyperbolic space.
@@ 1. Introduction
We point out here that the moving plane method on $\mathbb{H}^{n}$ was first seen in the work of Kumaresan-Prajapat [33, 34], where they established the analog result of Gidas-Ni-Nireneberg type, as well as solved an overdetermined problem on $\mathbb{H}^{n}$, which is an analogue of the problem on $\mathbb{R}^{n}$ initiated by Serrin [51] (see also [29], [48]). The hyperbolic space serves as one of the most important models of Riemannian manifolds with constant curvature. For definitions and basic properties of hyperbolic spaces, we refer to Section 2. The axial symmetry of solutions to a class of integral equations on half spaces was studied by Lu-Zhu in [49]. More recently, the overdetermined problem of fractional order equations on hyperbolic spaces was studied by Li-Lu-Wang [38]. The concept of the moving plane method in [33, 34] was further developed in the work of Almeida-Ge [4] and Almeida-Damascelli-Ge [3], where they took advantage of the foliation structure of $\mathbb{H}^{n}$ (for details see Section 2). Li, Lu and Yang [40] obtained a higher order symmetry result of the Brezis-Nirenberg problem on the hyperbolic spaces [40]. Recently. with Li and Lu, the author studied general higher order equations [37] by using the moving plane method, and later obtained the classification result using the moving sphere method [39] on the hyperbolic space. For symmetry and existence results of semilinear PDEs of second order on hyperbolic space, see Mancini and Sandeep [50], and [21] on a quasilienar equation involving the $p$-Laplacian on $\mathbb{H}^{n}$.
@@
$(-\Delta_{\mathbb{H}^{n}})^{s}u=u^{p}(x),\quad x\in\mathbb{H}^{n},$ (1.7)
where $1<p\leqslant p^{*}=\frac{n+2s}{n-2s}$. We say that a nonnegative function $u\in H^{s}(\mathbb{H}^{n})$ is a weak solution of (1.7) if for any test function $\varphi\in C_{c}^{\infty}(\mathbb{H}^{n})$, there holds
@@
We point out that the existence of nonnegative weak solutions has been showed recently by Bruno and Papageorgiou [10] for subcritical exponents. To be more precise, they considered the following more general problem on $\mathbb{H}^{n}$:
$(-\Delta_{\mathbb{H}^{n}})^{s}u-\lambda^{s}u-u^{p}=0,$
where $\lambda\in(0,1)$ and $1<p<p^{*}$, and showed that there exists at least one nontrivial nonnegative weak solution in $H^{s}(\mathbb{H}^{n})$. They also established the boundedness and regularity of the weak solutions, among many other results.
@@
Let $n\geqslant 2$, $0<s<1$. Suppose that $u\in H^{s}(\mathbb{H}^{n})$ is a nonnegative weak solution of (1.7). Then there exists a point $P\in\mathbb{H}^{n}$ such that $u$ is radially symmetric and nondecreasing about $P$.
@@
This equivalence plays an essential role in applying the moving plane method to prove Theorem 1.1. The Helgason-Fourier analysis on general symmetric spaces was first developed by Helgason [31, 32]. It has been effectively used in harmonic analysis and PDEs. In addition, the Helgason-Fourier analysis plays a key role in the establishment of higher order Hardy-Sobolev-Maz’ya’s inequalities and Green’s function of GJMS operators, which was studied by Lu and Yang [44, 45, 46] and Flynn, Lu, Yang [22]. Using Fourier transform, we are able to derive the precise expression of the Green’s function $G_{s}(\rho)$ and study its asymptotic behaviors near zero and infinity, which are the key ingredients in the moving plane method. For heat kernel and Green function estimates on general noncompact symmetric spaces, we refer to [5, 52], among many others.
@@
For $s\in(0,1)$, there holds that
@@ 2.1. Models of hyperbolic spaces
  • The hyperboloid model: It is given by the upper sheet of the two-sheeted hyperboloid in $\mathbb{R}^{n,1}$:
    $\mathbb{H}^{n}=\{x=(x_{0},x_{1},\cdots,x_{n})\in\mathbb{R}^{n,1}:x_{0}^{2}-x_{1}^{2}-\cdots-x_{n}^{2}=1,x_{0}>0\},$
    equipped with the metric induced from the Lorentzian metric on $\mathbb{R}^{n,1}$:
    $ds^{2}=-dx_{0}^{2}+dx_{1}^{2}+\cdots+dx_{n}^{2}.$
@@ 2.2. Hardy-Littlewood-Sobolev inequality
$\|I_{\alpha}f\|_{L^{q}(\mathbb{H}^{n})}\lesssim\|f\|_{L^{p}(\mathbb{H}^{n})}.$ (2.2)
@@ 2.3. Sobolev inequality for fractional Laplacians on $\mathbb{H}^{n}$
For a general survey and the manifold setting, one may be referred to Hebey [30].
@@ 2.4. The Helgason-Fourier transform on hyperbolic spaces
the generalized eigenfunctions of the Laplace-Beltrami operator $\Delta_{\mathbb{H}^{n}}$ on the ball model $\mathbb{B}^{n}$ with eigenvalue $(\lambda^{2}+\frac{(n-1)^{2}}{4})$. Equivalently, the eigenfunctions on the hyperboloid model is
$h_{\lambda,\theta}=[x,(1,\theta)]^{i\lambda-(n-1)/2}$
and satisfies $\Delta_{\mathbb{H}^{n}}h_{\lambda,\theta}=-(\lambda^{2}+\frac{(n-1)^{2}}{4})h_{\lambda,\theta}$. The Fourier transform of a function $f$ on $\mathbb{H}^{n}$ is defined as
@@
$f(x)=D_{n}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\hat{f}(\lambda,\zeta)e_{\lambda,\zeta}(x)|c(\lambda)|^{-2}d\lambda d\sigma,$
where $D_{n}=(2^{3-n}\pi|\mathbb{S}^{n-1}|)^{-1}$ and $c(\lambda)$ is the Harish-Chandra’s $c$-function given by
$c(\lambda)=\frac{2^{n-1-i\lambda}\Gamma(n/2)\Gamma(i\lambda)}{\Gamma(\frac{n-1+i\lambda}{2})\Gamma(\frac{1+i\lambda}{2})}.$
@@
$\int_{\mathbb{B}^{n}}|f(x)|^{2}dV=D_{n}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\hat{f}(\lambda,\zeta)|c(\lambda)|^{-2}d\lambda d\sigma.$
@@ 2.5. Foliations of hyperbolic spaces
Let $U=\mathbb{H}^{n}\cap\{x_{1}=0\}$ and $U_{t}=A_{t}(U)$, $\mathbb{H}^{n}$ is then foliated by $U_{t}$ and $\mathbb{H}^{n}=\bigcup_{t\in\mathbb{R}^{n}}U_{t}$. The reflection $I$ is an isometry such that $I^{2}=Id$ and $I$ fixes the hypersurface $U$, by $I(x_{0},x_{1},x_{2},\cdots,x_{n})=(x_{0},-x_{1},x_{2},\cdots,x_{n})$. Moreover, the reflection w.r.t. $U_{t}$ is defined as $I_{t}=A_{t}\circ I\circ A_{-t}$, and $U_{t}$ is fixed by $I_{t}$.
@@ 3. Fractional Laplacian on hyperbolic spaces
We point out here that the above two definitions of fractional Laplacians are not equivalent. Although the later is also defined via an extension problem, the manifold $\mathbb{H}^{n}\times(0,\infty)$ is not conformally compact Einstein. In fact, the former operator (Chang-González) lives on the conformal infinite of $X$ while the latter (Banica–González–Sáez) is defined on the bulk manifold $X$ itself. As a consequence, the fractional Laplacian we treat here does not coincide with the scattering operator on hyperbolic spaces, which is not conformally covariant.
@@
$L_{\lambda}(x,y)=\int_{\mathbb{S}^{n}}\bar{h}_{\lambda,\theta}(x)h_{\lambda,\theta}(y)\,d\theta.$ (3.2)
We note the following explicit formulae of $L_{s}(x,y)$ given by Banica [6].
@@
$L_{\lambda}(x,y)=C_{n}\,(\sinh\rho)^{-\frac{n-2}{2}}P_{-\frac{1}{2}+i\lambda}^{-\frac{n-2}{2}}(\cosh\rho),$ (3.3)
@@
$L_{\lambda}(\rho)=C_{n}\left(-\frac{1}{\sinh\rho}\frac{d}{d\rho}\right)^{\frac{n-1}{2}}\cos(\lambda\rho),$ (3.4)
@@
$L_{\lambda}(\rho)=C_{n}\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(-\frac{1}{\sinh r}\frac{d}{dr}\right)^{\frac{n}{2}}\cos(\lambda r)\,dr.$ (3.5)
The kernel of Green’s function $G_{s}(x,y)$ can be considered as the inverse of the fractional Laplacian. In [7], the authors provided the information of $\mathcal{K}_{s}$. The compuation for $G_{s}$ is similar, and we present the detailed proof here for completeness.
@@
$\Big(\frac{\partial_{\rho}}{\sinh\rho}\Big)^{m}\sim 2^{m}e^{-m\rho}\partial_{\rho}^{m}-(2^{m+1}e^{-m\rho}+2^{m}e^{-\rho})\partial_{\rho}^{m-1},\qquad\text{as }\rho\to\infty.$
Finally, we show the following monotonicity property of the kernel $G_{n,s}(\rho)$.
@@
We use the spectral representation of the Green kernel. Recall that
$G_{s}(\rho)=\int_{-\infty}^{\infty}\left(\lambda^{2}+\tfrac{(n-1)^{2}}{4}\right)^{-s}L_{\lambda}(\rho)\,|c(\lambda)|^{-2}\,d\lambda,$
where $L_{\lambda}(\rho)$ is given in Lemma 3.3. In particular, $L_{\lambda}(\rho)$ depends only on $\rho=d_{\mathbb{H}^{n}}(x,y)$ and is the spherical function associated with the eigenvalue $(\lambda^{2}+\tfrac{(n-1)^{2}}{4})$.
For radial function $f$, we have $\Delta_{\mathbb{H}^{n}}f(\rho)=f^{\prime\prime}(\rho)+(n-1)\coth(\rho)f^{\prime}(\rho)$. Thus,
$L_{\lambda}^{\prime\prime}(\rho)+(n-1)\coth(\rho)L_{\lambda}^{\prime}(\rho)=-(\lambda^{2}+\tfrac{(n-1)^{2}}{4})L_{\lambda}(\rho).$
Mulitplying both sides by $\sinh^{n-1}(\rho)$, we have
$(\sinh^{n-1}(\rho)L_{\lambda}^{\prime}(\rho))^{\prime}=-(\lambda^{2}+\tfrac{(n-1)^{2}}{4})\sinh^{n-1}(\rho)L_{\lambda}(\rho)<0.$
Thus, $\sinh^{n-1}(\rho)L_{\lambda}^{\prime}(\rho)$ is strictly decreasing in $\rho$, which implies $L_{\lambda}^{\prime}(\rho)<0$ and $G_{s}^{\prime}(\rho)<0$. ∎
@@
Lemma 4.1.
Let $s\in(0,1)$ and $1<p\leqslant\frac{n+2s}{n-2s}$. Assume $u\geq 0$ and
$u\in L^{\frac{2n}{n-2s}}(\mathbb{H}^{n}).$ (4.1)
If $p<\frac{n+2s}{n-2s}$, assume in addition that
$u^{p}\in L^{\frac{2n}{n+2s}}(\mathbb{H}^{n}).$ (4.2)
Then if $u$ is a weak solution of
$(-\Delta_{\mathbb{H}^{n}})^{s}u=u^{p}\quad\text{in }\mathbb{H}^{n},$ (4.3)
then $u$ satisfies the integral equation
$u(x)=\int_{\mathbb{H}^{n}}G_{s}\big(\rho(x,y)\big)\,u(y)^{p}\,dV_{y}\qquad\text{for a.e. }x\in\mathbb{H}^{n},$ (4.4)
where $G_{s}$ is the Green’s function defined in (1.9). The converse also holds, that is, if $u$ satisfies (4.4), then $u$ is a weak solution of (4.3).
@@
$(-\Delta_{\mathbb{H}^{n}})^{s}\varphi(x)=\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\Bigl(\lambda^{2}+\frac{(n-1)^{2}}{4}\Bigr)^{s}\widehat{\varphi}(\lambda,\theta)\,\bar{h}_{\lambda,\theta}(x)\,|c(\lambda)|^{-2}\,d\theta\,d\lambda.$ (4.5)
@@
$\int_{\mathbb{H}^{n}}u(x)(-\Delta_{\mathbb{H}^{n}})^{s}\varphi(x)dV_{x}$ $=\int_{\mathbb{H}^{n}}u(x)^{p}\varphi(x)dx.$
@@
$\int_{\mathbb{H}^{n}}u(x)(-\Delta_{\mathbb{H}^{n}})^{s}\varphi(x)\,dx$
$=\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\Bigl(\lambda^{2}+\tfrac{(n-1)^{2}}{4}\Bigr)^{s}\widehat{\varphi}(\lambda,\theta)\left[\int_{\mathbb{H}^{n}}u(x)\bar{h}_{\lambda,\theta}(x)\,dx\right]|c(\lambda)|^{-2}\,d\theta\,d\lambda.$
Using the inversion formula, the right side becomes
$\int_{\mathbb{H}^{n}}u(x)^{p}\varphi(x)\,dx$ $=\int_{\mathbb{H}^{n}}u(x)^{p}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\widehat{\varphi}(\lambda,\theta)\bar{h}_{\lambda,\theta}(x)|c(\lambda)|^{-2}\,d\theta\,d\lambda\,dV_{x}$
$=\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\widehat{\varphi}(\lambda,\theta)\left[\int_{\mathbb{H}^{n}}u(x)^{p}\bar{h}_{\lambda,\theta}(x)\,dV_{x}\right]|c(\lambda)|^{-2}\,d\theta\,d\lambda.$
@@
$\Bigl(\lambda^{2}+\tfrac{(n-1)^{2}}{4}\Bigr)^{s}\int_{\mathbb{H}^{n}}u(x)\bar{h}_{\lambda,\theta}(x)\,dV_{x}=\int_{\mathbb{H}^{n}}u(x)^{p}\bar{h}_{\lambda,\theta}(x)\,dV_{x}$
Equivalently,
$(\mu^{2}+\rho_{0}^{2})^{s}\,\widehat{u}(\mu,\theta)=\widehat{u^{p}}(\mu,\theta).$
Thus, by inversion formula again, we obtain
$u(x)$ $=\int_{\mathbb{H}^{n}}u(y)^{p}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}\left(\lambda^{2}+\tfrac{(n-1)^{2}}{4}\right)^{-s}\bar{h}_{\lambda,\theta}(x)h_{\lambda,\theta}(y)|c(\lambda)|^{-2}d\theta d\lambda dV_{y}$